# Dual-Personalizing Adapter for Federated Foundation Models

### v1

@@ 7 Conclusion
In this work, we propose a novel setting in federated foundation models by considering the test-time distribution shifts. In this setting, it is imperative for models to not only excel in performance on targeted local datasets but also to yield comparable results in test-time tasks. To address these challenges, we present a new method FedDPA, consisting of two core components: a global adapter to acquire general knowledge for test-time tasks, and a local adapter to learn personalized information for targeted local tasks. To further enhance the performance, an instance-wise dynamic weighting mechanism is introduced for balancing the local and global adapters during the inference. Compared with existing methods, our method achieves promising results on the constructed federated datasets with various NLP tasks, demonstrating its effectiveness.
@@ A.1 Datasets
In this paper, we have developed two federated datasets derived from the Flan Wei et al. [2021], and details to construct these datasets are elucidated in this section. Flan encompasses a diverse array of NLP tasks, each comprising multiple datasets. To align with FL settings, we employed a stratified selection process, randomly choosing one dataset from each of the eight distinct tasks from Flan to form each federated dataset. In addition, to simulate client local data scarcity McMahan et al. [2017], we implemented a downsampling strategy, reducing the size of each selected dataset to 300 training instances and 200 testing instances. Consequently, each constructed federated dataset encompasses eight distinct NLP tasks, with a total of 300 training examples and 200 testing examples per task, culminating in a whole dataset comprising 2400 training examples and 1600 testing examples across all tasks. The specific tasks and datasets included in each federated dataset are cataloged in Table 5.
@@
In our ablation study in section 6.2 examining the client sample rate to align with FL settings, we divided each task in our constructed federated datasets into five subsets, each comprising an equal number of training data. Based on our assumption that each client is associated with a single task, this division results in a total of 40 clients, with each client possessing a dataset of 60 training examples. To mimic real-world FL communication dynamics, we employed a randomized selection process for clients (subsets) within each task according to specified sample rates. Accordingly, for sample rates specified as $\{0.2,0.4,0.6,0.8,1\}$, we selected 1,2,3,4, and 5 clients (subsets) per task, leading to 8, 16, 24, 32, and 40 clients participating in federated communications, respectively. The evaluation phase involves computing the average results across these selected clients for each specified sample rate, which provides a comprehensive analysis of how client sample rates influence the performance of our method.
@@
Input: The father convinced his son that it is possible for him to one day become a knight, but he may never achieve such status coming from a peasant family. Who is ”he”? OPTIONS: - The father - his son
@@
Table 6: Examples of data in our constructed federated datasets.
@@
Table 7: Prompt Template.
A.2 Baselines
In this section, detailed descriptions of the implementation for each baseline compared in this study will be provided:
@@ A.2 Baselines
  • Local-finetuned model: This model trains independently of any external communication with other clients or a central server. It is specifically trained on data pertaining to a single task, dedicating 50 epochs to optimize for task-specific performance without the influence of external data.
@@
  • FedIT Zhang et al. [2023a]: The FedIT model is the final aggregated global model derived from diverse local client datasets after training. It embodies the essence of collaborative learning inherent to federated learning, assimilating knowledge from a multitude of client-specific data sources.
@@
  • FedLoRA model Yi et al. [2023]: Here, we adapt the training paradigm in paper Yi et al. [2023] to NLP tasks by focusing on training the lightweight LoRA for aggregation while keeping the majority of the LLM parameters frozen. Subsequently, a personalized adaptation process is employed, where the globally aggregated LoRA undergoes further refinement on each local client’s dataset to tailor the learning outcomes to individual client needs.
@@ Appendix B Additional Experiments
B.1 Instance-Wise Dynamic Weighting Mechanism Analysis
In this section, we further examine the impact of the instance-wise dynamic weighting mechanism, including the selected local instance number and the type of instance representation.
@@ B.1 Instance-Wise Dynamic Weighting Mechanism Analysis
In Section 4.4, our method entails utilizing the embedding of the final token from the last hidden layer of the LLM, denoted as ’LAST’, as the input instance representation for the purpose of similarity calculation. In this exploration, we delve into another instance representation strategy, which involves employing the average embedding of all tokens from the final hidden layer of the LLM, herein referred to as ’AVG’. The comparative analysis, as presented in Table 8, demonstrates that employing the embedding of the last token yields superior performance relative to the strategy of averaging the embeddings of all tokens. This observed difference in performance can be attributed to the decoder structure inherent to LLMs, wherein the final token is capable of attending to all preceding tokens, thereby encapsulating comprehensive sentence-level information.
Methods Emb Fed Dataset 1
P TTP
FedDPA-F AVG 79.30 76.77
LAST 79.48 77.12
FedDPA-T AVG 79.65 73.36
LAST 80.22 75.47
Table 8: Ablation study of instance representations (Emb). P represents personalization, and TTP represents test-time personalization. LAST represents using the embedding of the final token from the final hidden layer of LLM as instance representation, and AVG represents using the average embedding of all tokens from the final hidden layer of LLM as instance representation.
B.2 Model Scalability Analysis
@@
B.3 Communication and Computation Analysis

### v2

@@ 7 Conclusion
Federated Foundation Model (FedFM) is a promising direction to enhance existing Foundation Models, e.g. LLM, by leveraging private data sources. Test-time distribution shift is a critically important problem to ensure the practicability of the FedFM system. This work is the first to propose the test-time FedFM setting. To tackle this challenging scenario, we propose a novel dual-personalizing adapter for the FedFM framework. The method is evaluated on public NLP tasks that are adapted to mimic the test-time FedFM setting. This work is the first step towards this direction. We focus on defining a new learning scenario, proposing a basic learning framework, and setting up the benchmark datasets. Our future works will be in two directions: the first is to rethink this problem from a theoretical perspective, and the second is to enhance the benchmark setting with more datasets in real applications.
@@ A.1 Datasets
In this paper, we have developed two federated datasets derived from the Flan [31], and details to construct these datasets are elucidated in this section. Flan encompasses a diverse array of NLP tasks, each comprising multiple datasets. These tasks, generated from different contextual factors, inherently experience various complex distribution shifts. To align with FL settings, we employed a stratified selection process, randomly choosing one dataset from each of the eight distinct tasks from Flan to form each federated dataset. In addition, to simulate client local data scarcity [24], we implemented a downsampling strategy, reducing the size of each selected local dataset to 300 training instances and 200 testing instances. Consequently, each constructed federated dataset encompasses eight distinct NLP tasks, culminating in a whole dataset comprising 2400 training examples and 1600 testing examples across all tasks. The specific tasks and datasets included in each federated dataset are cataloged in Table 5.
@@
In our ablation study in section 6.2 examining the client number to align with FL settings, we divided each task in our constructed federated datasets into five subsets, each comprising an equal number of training data. Based on our assumption that each client is associated with a single task, this division results in a total of 40 clients, with each client possessing a local dataset of 60 training examples. To mimic real-world FL communication dynamics, we employed a randomized selection process for clients (subsets) within each task according to specified sample rates. Accordingly, for sample rates specified as $\{0.2,0.4,0.6,0.8,1\}$, we selected 1,2,3,4, and 5 clients (subsets) per task, leading to 8, 16, 24, 32, and 40 clients participating in federated communications, respectively. The evaluation phase involves computing the average results across these selected clients for each specified sample rate, which provides a comprehensive analysis of how client numbers influence the performance of our method.
Different with Multi-Task Learning.
Although both multi-task learning and our test-time personalization setting involve multiple tasks during training and testing, multi-task learning operates in a centralized setting, whereas our setting is based on federated learning, a distributed setting. More importantly, our setting accounts for test-time distribution shifts, a challenge that is typically overlooked in conventional multi-task learning. Additionally, fine-tuning on the combination of multi-task data from a centralized foundation model serves as a strong baseline for multi-task learning. In foundation models, all tasks are standardized into a uniform format, and the model benefits from task-agnostic token embeddings learned through extensive pre-training on diverse data. Thus, directly fine-tuning on this multi-task data represents the implementation of multi-task learning using foundation models  [40]. We have included this baseline, referred to as "Centralized," in our experimental comparisons.
@@
Table 6: Examples of data in our constructed federated datasets.
@@
Input: The father convinced his son that it is possible for him to one day become a knight, but he may never achieve such status coming from a peasant family. Who is "he"? OPTIONS: - The father - his son
@@
Table 7: Prompt Template.
@@
A.2 Baselines and Implementation
In this section, detailed descriptions of the implementation of FedDPA and each baseline compared in this study will be provided:
@@ A.2 Baselines and Implementation
  • Local-finetuned model: This model trains independently without any external communication from other clients or a central server. It is specifically trained on data pertaining to a single task, dedicating 50 epochs to optimize for task-specific performance without the influence of external data.
@@
  • FedIT model [42]: The FedIT model is the final aggregated global model derived from diverse local client datasets after training. It embodies the essence of collaborative learning inherent to federated learning, assimilating knowledge from a multitude of client-specific data sources.
@@
  • FedLoRA model [38]: Here, we adapt the training paradigm in paper [38] to NLP tasks by focusing on training the lightweight LoRA for aggregation while keeping the majority of the LLM parameters frozen. Subsequently, a personalized adaptation process is employed, where the globally aggregated LoRA undergoes further local training on each local client’s dataset to tailor the learning outcomes to individual client needs.
  • FedDPA-F: FedDPA-F is the combination of the global adapter and the local fine-tuning adapter. During the inference, the scale factor is set to $\lambda=1$ in the instance-wise dynamic weighting mechanism.
  • FedDPA-T: FedDPA-T is the combination of the global adapter and the local training adapter. Since the global adapter contributes to the training of the local adapter for personalization, it is essential to restrict the similarity score during the inference. This adjustment is necessary to ensure that the integration of the global and local adapters achieves optimal personalization outcomes. Thus, the scale factor $\lambda$ is set equal to the updating weight $\alpha$ used in the local adapter training.
All models are implemented using LoRA to enhance learning efficiency, with the rank of LoRA set as $r=8$ and only applied to ${\bm{W}}_{q}$ and ${\bm{W}}_{v}$. For FL methods, each client conducts 10 local epochs with a batch size of 32. We implement all the methods using PyTorch and conduct all experiments on NVIDIA Quadro RTX 6000 GPU.
A.3 Algorithm
Require : Number of clients $M$; number of communication rounds $K$; local step size $\eta$; freeze foundation model ${\bm{\theta}}$; initial global adapter $\Delta{{\bm{\theta}}}_{g}$ and local adapters ${\Delta{{\bm{\theta}}}_{l}^{1},\cdots,\Delta{{\bm{\theta}}}_{l}^{M}}$; local training datasets ${\bm{D}}^{1}_{train},\cdots,{\bm{D}}^{M}_{train}$.
// Learn Global Adapter
for $k\leftarrow 1$ to $K$ do
       Sample clients $\mathcal{S}\subseteq\{1,\cdots,M\}$;
       Communicate $\Delta{{\bm{\theta}}}_{g}$ to all clients $m\in\mathcal{S}$;
       for each client $m\in\mathcal{S}$ in parallel do
             Initialize $\Delta{{\bm{\theta}}}_{g}^{m}\leftarrow\Delta{{\bm{\theta}}}_{g}$;
             $\Delta{{\bm{\theta}}}_{g}^{m}\leftarrow\textbf{Client local training}(\{{\bm{\theta}},\Delta{{\bm{\theta}}}_{g}\},{\bm{D}}^{m}_{train},\eta)$; [Equation 2]
             Communicate $\Delta{{\bm{\theta}}}_{g}^{m}$ to the server;
      Construct $\Delta{{\bm{\theta}}}_{g}=\sum_{m\in\mathcal{S}}\frac{1}{|\mathcal{S}|}\Delta{{\bm{\theta}}}_{g}^{m}$; [Equation 2]
// Learn Local Adapters
for each client $m\in\{1,\cdots,M\}$ in parallel do
       Initialize $\Delta{{\bm{\theta}}}_{l}^{m}\leftarrow\Delta{{\bm{\theta}}}_{g}$;
       $\Delta{{\bm{\theta}}}_{l}^{m}\leftarrow\textbf{Client local training}(\{{\bm{\theta}},\Delta{{\bm{\theta}}}_{l}^{m}\},{\bm{D}}^{m}_{train},\eta)$; [Equation 3]
return $\Delta{{\bm{\theta}}}_{g},\{\Delta{{\bm{\theta}}}_{l}^{1},\cdots,\Delta{{\bm{\theta}}}_{l}^{M}\}$.
Algorithm 1 FedDPA-F
Require : Number of clients $M$; number of communication rounds $K$; local step size $\eta$; freeze foundation model ${\bm{\theta}}$; initial global adapter $\Delta{{\bm{\theta}}}_{g}$ and local adapters ${\Delta{{\bm{\theta}}}_{l}^{1},\cdots,\Delta{{\bm{\theta}}}_{l}^{M}}$; local training datasets ${\bm{D}}^{1}_{train},\cdots,{\bm{D}}^{M}_{train}$.
for $k\leftarrow 1$ to $K$ do
       Sample clients $\mathcal{S}\subseteq\{1,\cdots,M\}$;
       Communicate $\Delta{{\bm{\theta}}}_{g}$ to all clients $m\in\mathcal{S}$;
       for each client $m\in\mathcal{S}$ in parallel do
             Initialize $\Delta{{\bm{\theta}}}_{g}^{m}\leftarrow\Delta{{\bm{\theta}}}_{g}$ and $\Delta{{\bm{\theta}}}_{l}^{m}\leftarrow\Delta{{\bm{\theta}}}_{l}^{m,k-1}$;
             // Learn Global Adapter
             $\Delta{{\bm{\theta}}}_{g}^{m}\leftarrow\textbf{Client local training}(\{{\bm{\theta}},\Delta{{\bm{\theta}}}_{g}\},{\bm{D}}^{m}_{train},\eta)$; [Equation 2]
             // Learn Local Adapter
             $\Delta{{\bm{\theta}}}_{l}^{m,k}\leftarrow\textbf{Client local training}(\{{\bm{\theta}},\Delta{{\bm{\theta}}}_{g},\Delta{{\bm{\theta}}}_{l}^{m}\},{\bm{D}}^{m}_{train},\eta)$; [Equation 3]
             Communicate $\Delta{{\bm{\theta}}}_{g}^{m}$ to the server;
             Maintain $\Delta{{\bm{\theta}}}_{l}^{m,k}$ locally;
      Construct $\Delta{{\bm{\theta}}}_{g}=\sum_{m\in\mathcal{S}}\frac{1}{|\mathcal{S}|}\Delta{{\bm{\theta}}}_{g}^{m}$; [Equation 2]
return $\Delta{{\bm{\theta}}}_{g},\{\Delta{{\bm{\theta}}}_{l}^{1},\cdots,\Delta{{\bm{\theta}}}_{l}^{M}\}$.
Algorithm 2 FedDPA-T
@@ Appendix B Additional Experiments
B.1 Adaptability Analysis
To enhance applicability across diverse non-IID environments, our method is meticulously designed with a high degree of flexibility for its adaption across various global learning frameworks, backbones and PEFT methods for different scenarios. This adaptability is simply achieved through the straightforward substitution of the FedAvg, LLM and LoRA with alternative aggregation methods, transformer-based foundation models and adapter-based PEFT methods during the training. In our experiment, we employ FedAvg, LLM and LoRA as representative examples, demonstrating our methods’ superior performance compared to other baselines as indicated in Table 1 and Table 2. To further validate the effectiveness and versatility of our approach within different federated learning contexts, we adapt our methods to include the FedProx[20] framework and also implement other baselines (FedIT and FedLoRA) within FedProx to maintain a fair comparison.
Results presented in Table 8 indicate that our methods, both training and fine-tuning methods, outperform competing approaches. Specifically, FedDPA-T excels in personalization, while FedDPA-F leads in test-time personalization, maintaining consistent performance with FedAvg. These findings underscore the robustness and consistent efficacy of our methods across various global learning paradigms for different non-IID scenarios.
Table 8: Personalization and test-time personalization results of different models with FedProx framework on federated dataset 1. FedDPA-F represents the model with the local fine-tuning adapter and FedDPA-T represents the model with the local training adapter. Linguistic represents the linguistic acceptability task, Word Dis represents the word disambiguation task, and Question CLS represents question classification task.
Methods Federated Dataset 1
Personalization
FedIT 71.00 83.00 70.38 96.32 69.50 63.50 73.55 91.50 77.34
FedLoRA 79.00 85.00 70.63 96.60 69.00 63.00 78.85 88.50 78.82
FedDPA-F 77.50 85.00 71.92 96.78 73.00 63.50 77.87 89.50 79.38
FedDPA-T 79.50 85.50 71.62 96.89 76.00 60.00 77.28 93.50 80.04
Test-Time Personalization
FedLoRA 76.76 75.82 74.11 74.36 75.01 72.27 77.16 74.87 75.05
FedDPA-F 77.25 76.01 76.72 76.95 77.21 75.48 77.31 76.14 76.63
FedDPA-T 75.64 74.62 75.58 75.01 74.90 74.21 74.80 75.35 75.01
B.2 Instance-Wise Dynamic Weighting Mechanism Analysis
In this section, we further examine the impact of the instance-wise dynamic weighting mechanism, including the similarity metric, the selected local instance number and the type of instance representation.
Impact of Similarity Metric.
In Section 4.3, we employ the similarity metric to calculate the average similarity scores of each input instance, which serves as the weight $\alpha_{t}$ to dynamically balance the global and local adapters. For this purpose, cosine similarity is selected in our experiment due to its better robustness and normalization with high-dimensional vectors than other metrics, and its superiority has been demonstrated in many NLP/CV works[25, 37, 36]. Additionally, we conducted an ablation study comparing other metrics like L2-norm and Pearson correlation, and the results in Table B.2 demonstrate that cosine similarity outperforms other similarity metrics.
@@ B.2 Instance-Wise Dynamic Weighting Mechanism Analysis
In Section 4.3, our method entails utilizing the embedding of the final token from the last hidden layer of the LLM, denoted as ’LAST’, as the input instance representation for the purpose of similarity calculation. In this exploration, we delve into another instance representation strategy, which involves employing the average embedding of all tokens from the final hidden layer of the LLM, herein referred to as ’AVG’. The comparative analysis, as presented in Table B.2, demonstrates that employing the embedding of the last token yields superior performance relative to the strategy of averaging the embeddings of all tokens. This observed difference in performance can be attributed to the decoder structure inherent to LLMs, wherein the final token is capable of attending to all preceding tokens, thereby encapsulating comprehensive sentence-level information. Table 9: Ablation study of similarity metric (Sim). P represents personalization, and TTP represents test-time personalization. -L2 represents using the L2-Norm as metric, Pearson represents using the Pearson correlation as metric, and Cosine represents using the cosine similarity as the metric. Methods Sim Fed Dataset 1 P TTP FedDPA-F -L2 77.69 77.10 Pearson 79.09 76.78 Cosine 79.48 77.12 FedDPA-T -L2 77.58 77.30 Pearson 79.73 75.27 Cosine 80.22 75.47 Table 10: Ablation study of instance representations (Emb). P represents personalization, and TTP represents test-time personalization. LAST represents using the embedding of the final token from the final hidden layer of LLM as instance representation, and AVG represents using the average embedding of all tokens from the final hidden layer of LLM as instance representation. Methods Emb Fed Dataset 1 P TTP FedDPA-F AVG 79.30 76.77 LAST 79.48 77.12 FedDPA-T AVG 79.65 73.36 LAST 80.22 75.47
B.3 Model Scalability Analysis
@@ B.4 Communication and Computation Analysis
Appendix C Discussions
C.1 Extend to Other Scenarios of Test-Time FL
In this paper, we primarily consider an ideal scenario for our proposed test-time personalization setting, where all tasks are included for all clients. In this section, we will discuss our methods under alternative scenarios.
In practical applications, it is possible that some tasks remain unseen by all clients during training and may only appear during the testing phase for certain clients. In this scenario, test-time distribution shifts arise from these unseen tasks. According to previous works [5, 29], generic features learned through FL are robust to distribution shifts, even those originating from unseen test-time tasks. Consequently, our method can be directly adapted to this scenario. We conducted experiments to evaluate our methods against other baselines on three unseen test-time tasks. All methods and baselines are trained on our constructed Federated Dataset 1 and tested on three tasks not included in Federated Dataset 1. For personalized methods, we report the average score across all clients and the maximum score among all clients to provide a comprehensive comparison.
As shown in Table 14, FedDPA-T outperforms all other models, indicating the effectiveness of our method on unseen test-time tasks. Additionally, FedDPA-F surpasses FedLoRA, suggesting that the generic features learned through FL across diverse data distributions are robust to various distribution shifts, consistent with the findings in [29]. Despite this, other techniques targeting these unseen test-time tasks could further enhance our proposed methods, which we leave for future work to explore.
Table 14: Test-time performance on unseen tasks. All models are trained on Federated Dataset 1, and these unseen test-time tasks are not included in Federated Dataset 1. AVG represents the average score across all clients for this task, while MAX represents the highest score among these clients for this task. The best performance for AVG is bolded, and the best performance for MAX is underlined.
Test-Time Task FedIT FedLoRA FedDPA-F FedDPA-T
AVG MAX AVG MAX AVG MAX
Summarization 22.40 22.25 22.42 22.37 22.52 22.46 22.65
Reading Comprehension 69.50 68.88 73.00 69.44 72.00 71.88 76.00
Open Domain QA 78.32 76.46 78.96 78.01 79.61 78.76 79.92
Our methods are also applicable to scenarios involving the introduction of new clients. As previously analyzed, our methods are robust to different distribution shifts. By computing the similarity between instances from new clients and existing clients through our instance-wise dynamic weighting mechanism, we can identify the most similar existing client. The model of this identified client can then be used as the initial model for the new clients, providing a more effective starting point for further training.
C.2 Limitations
The proposed dual-personalizing adapter architecture is limited by 1) model scales: the proposed methods rely on the foundation model, presupposing that each client possesses sufficient memory capacity and computational resources to store and train the foundation model with PEFT methods. 2) secure issues: the framework operates under the assumption that all clients are trusted and legally entitled to access and utilize data stored on them, and the whole process does not suffer from any attacks.